OpenStudy (anonymous):

-2/3(x - 1) = 36. Anyone know the answer? Please help!

OpenStudy (anonymous):

-2/3 (x-1)= 36 x-1 = 36 (-3/2) x-1 = 18(-3) x-1 = -54 x = -53

OpenStudy (anonymous):

Thank you so much, jerwyn gayo!!